OpenStudy (calleyleeann):

Find the derivative of f(x) = 5x + 9 at x = 2. I know this question has already been asked, but I don't understand the other answers.

OpenStudy (calleyleeann):

Using the difference quotient, I get 0, but other sources tell me the answer is 5. How do I get the right answer using the difference quotient?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

for a line, derivative = slope

OpenStudy (misty1212):

we can use the :"difference quotient" if you like, it is the change in y over the change in x , which is still the slope you need the steps?

OpenStudy (calleyleeann):

Let's use the difference quotient please :) And yes, let's go over the steps because I'm obviously doing something wrong with it lol

OpenStudy (misty1212):

ok first of all, the "at x = 2" part is unnecessary because the slope of a line is a constant, and therefore does not depend on the number but since that is what it says, lets go ahead and use it are you using \[\frac{f(x)-f(a)}{x-a}\] or \[\frac{f(a+h)-f(a)}{h}\]?

OpenStudy (calleyleeann):

Number 2!

OpenStudy (misty1212):

ok so first replacing \(x\) by \(2\) we have \[\frac{f(2+h)-f(2)}{h}\]

OpenStudy (misty1212):

in this case \[f(x)=5x+9\] so \[f(2)=5\times 2+9=19\]

OpenStudy (calleyleeann):

When plugging it in isn't it supposed to look like this?:\[\frac{ (5(x+h)+9)-(5x+9) }{ h }\]

OpenStudy (misty1212):

it is if you use x instead of 2

OpenStudy (misty1212):

since it makes no difference lets do it your way

OpenStudy (calleyleeann):

Oh, okay. I didn't realize you could plug in 2 right away - I just always did it at the end

OpenStudy (misty1212):

lets do it your way anyway it is pretty easy since all you have to do is some algebra in the numerator multiply out and combine like terms

OpenStudy (misty1212):

you can almost do it in your head \[5x+5h+9-5x-9=5h\]

OpenStudy (calleyleeann):

Wow, I just realized my mistake lol.I forgot to divide by h at the end and that totally threw me off! Thanks so much for your help :D

OpenStudy (misty1212):

lol your welcome \[\color\magenta\heartsuit\]